1.2H 2H I 11,1 2H 2H   2.   2H 2H 2H 2H 1H


I am trying to learn how to interpret 1H-NMR for the two above spectrums. How would you interpret these 1H-NMR spectrums to determine the shape of the molecule whose formula is C13H9NO3?
Answer #1

Formula of the compound is C13H9NO3.

Double bond equivalence = C + 1 - H/2 + N/2 = 13 +1 - 9/2 + 1/2 = 10

Stucture of the compound is:

Analysis of 1H NMR:

2Ha ; doublet; 8.35 ppm

2Hb; doublet; 7.95 ppm

2 Hc; doublet; 7.8 ppm

1He; triplet; 7.65 ppm

2 Hd; triplet; 7.50 ppm